# Mixed Precision Iterative Refinement for Linear Inverse Problems

### v1

@@ Abstract
This study investigates the iterative refinement method applied to the solution of linear discrete inverse problems by considering its application to the Tikhonov problem in mixed precision. Previous works on mixed precision iterative refinement methods for the solution of symmetric positive definite linear systems and least-squares problems have shown regularization to be a key requirement when computing low precision factorizations. For problems that are naturally severely ill-posed, we formulate the iterates of iterative refinement in mixed precision as a filtered solution using the preconditioned Landweber method with a Tikhonov-type preconditioner. Through numerical examples simulating various mixed precision choices, we showcase the filtering properties of the method and the achievement of comparable or superior accuracy compared to results computed in double precision as well as another approximate method.
@@
mixed precision, iterative refinement, Tikhonov regularization, preconditioned iterative methods, Landweber
@@ 1 Introduction
In response to advancements in GPU computing architectures that natively support mixed precision computing, Higham and Pranesh in [18] consider the solution of symmetric positive definite (SPD) linear systems and least-squares problems in multiple precisions using iterative refinement (IR). They found that definiteness of SPD matrices could be lost when rounded to a precision with fewer representation bits, which in turn, could result in the failure of a Cholesky factorization when adapting GMRES-based IR strategies; see [2, 3]. It was determined that the definiteness of the matrix could be preserved by shifting the diagonal either before or after rounding - which may be viewed as a form of regularization. Our interests in this work are focused on applying IR to least-squares problems whose matrix $A$ is naturally severely ill-conditioned without any assumed underlying structure, to which Tikhonov regularization is often applied. Other works by Higham and others concerning the use of IR to solve linear systems using mixed precision include [12, 19, 23, 4].
@@ 2.1 Filtered Solution of Linear Discrete Ill-posed Problems
Many inverse problems whose approximate solution is given by (1) have been well studied using the singular value decomposition (SVD) of $A=U\Sigma V^{T}$, where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ are orthogonal matrices whose columns represent the left and right singular vectors of $A$, respectively. The diagonal matrix $\Sigma\in\mathbb{R}^{m\times n}$ contains the singular values on its main diagonal ordered in non-increasing fashion. Using the SVD, one can write the least-squares solution of (1) as a linear combination of the right singular vectors with the $j^{th}$ coefficient given by $u_{j}^{T}b/\sigma_{j}$. Since the singular values of $A$ cluster near numerical zero, the error $e$ contaminating $b$ is propagated into the computed solution for high index values $j$ corresponding to high frequency information [15, 14].
@@ 2.2 Landweber Iteration
where $\zeta\in(0,2\sigma_{1}^{2})$ is a relaxation term with $\sigma_{1}$ the largest singular value of $A$. Without loss of generality, we choose $\zeta=1$ in this work since $A$ can be scaled appropriately. While robust, this classical variant of Landweber is well known to have slow convergence towards a useful solution which can be understood from viewing its $k^{th}$ solution as a filtering method [13]:
(7) $x^{(k)}=\sum_{j=1}^{n}\left(1-(1-\sigma_{j}^{2})^{k}\right)\frac{u_{j}^{T}b}{\sigma_{j}}v_{j}$
@@ 2.3 Iterative Refinement on the Tikhonov Problem
Variants of fixed and mixed precision iterative refinement (IR) methods are not new - having been investigated consistently since Wilkinson’s first programmed version in 1948 [31]. A historical treatise of the progression of IR for linear problems may be found in [2, 16]. To compute the $k^{th}$ approximate solution, $x^{(k)}$, to the available linear system the $(k-1)^{th}$ residual $r^{(k-1)}=b-Ax^{(k-1)}$ is first computed. Then the correction equation $Ah=r^{(k-1)}$ is solved and the current solution is updated $x^{(k)}=x^{(k-1)}+h^{(k-1)}$. The process may be repeated as necessary until a termination criterion is met. When considering the application of IR to the Tikhonov problem given by
@@
which was first discussed by Riley in [27], Golub in [11] noted that the IR procedure is equivalent to iterated Tikhonov regularization whose $(k+1)^{th}$ iterate is given by
(9) $x^{(k+1)}=x^{(k)}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}r^{(k)}.$
@@
1  Input: $A\in\mathbb{R}^{m\times n},\,\,b\in\mathbb{R}^{m},\,\,x^{(0)}\in\mathbb{R}^{n},\,\,\alpha>0$
2  Output: $x^{(k+1)}\in\mathbb{R}^{n}$
@@ 7        $x^{(k+1)}=x^{(k)}+h^{(k)}$
From Algorithm 1, the $k^{th}$ iterate of IR on the Tikhonov problem may be written recursively as
$x^{(k)}=x^{(k-1)}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(A^{T}A+\alpha^{2}I\right)^{-1}x^{(k-1)}.$
Here, each iterate involves the computation of a Tikhonov regularized linear system with $A$ and $r^{(k-1)}$. The $k^{th}$ filter factors take the form
@@
which are described in [11] and [20]. Differently than discussed in the aforementioned works, we show that the filter factors of IR on the Tikhonov problem can be written with respect to the $k^{th}$ iterate of iterated Tikhonov regularization (9). This framework lends itself to the filter factor analysis in mixed precision where instead of the $k^{th}$ iterate of (9) we utilize the $k^{th}$ iterate of preconditioned Landweber with a Tikhonov-type preconditioner, which we show in the next section can be written as a filtered solution.
@@ 3 Landweber with a Tikhonov-type Preconditioner
(12) $\hat{x}^{(k+1)}=x^{(k)}+\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax^{(k)}\right).$
@@
Herein, we utilize the $M$ notation to help simplify our notation throughout this work and to consider the situation that $M^{T}M$ represents an approximate Tikhonov-type preconditioner to $A^{T}A+\alpha^{2}I$. Possible approximations we consider in Section 5 include that of (1) $M^{T}M$ as a low precision SVD approximation and (2) when the structure of $A$ allows for the approximation by a second matrix $C$ whose structure allows for fast and robust matrix-vector products. The former will be our focus when considering IR on the Tikhonov problem in mixed precision. The later case can arise naturally, for example, in image deblurring problems; see Section 5.
@@
To investigate the filter factors of mixed precision IR on the Tikhonov problem we first describe in Proposition 1 the filter factors of preconditioned Landweber with $M^{T}M=\left(A^{T}A+\alpha^{2}I\right)$. We again emphasize that the main use of the $M$ notation is to delineate between the possible change in structure of the preconditioner $M$ or its representation in a lower precision. Throughout, we will use a capital letter subscript, Q, to denote terms that come from a representation of $Q$, e.g. the SVD of $Q$. Additionally, we utilize subscripts $PL$ and $IR$ on vectors to denote solutions from preconditioned Landweber and IR, respectively.
@@
(14) $x^{(k)}_{PL}=\sum_{j=1}^{n}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{k}\right)\frac{u_{A,j}^{T}b}{\sigma_{j}}v_{M,j}$
@@
where we recall that $\hat{A}=AM^{-1}$ and $\hat{x}=Mx$. Using our definition of $M$ as $M=D_{M}V_{M}^{T}$ and the SVD of $A$, we may rewrite $\hat{A}$ as
@@
(17) $\begin{split}x^{(k)}_{PL}&=M^{-1}\hat{x}^{(k)}_{PL}\\ &=V_{M}D_{M}^{-1}\hat{D}^{k-1}D_{M}^{-1}\Sigma_{A}^{T}U_{A}^{T}b\\ &=\sum_{j=1}^{n}\hat{d}_{j}^{k-1}\frac{\sigma_{j}}{d_{M,j}^{2}}\left(u_{A,j}^{T}b\right)v_{M,j}\end{split}$
@@
$\begin{split}\hat{d}_{j}^{k-1}&=\sum_{i=0}^{k-1}\left(1-\frac{\sigma_{j}^{2}}{d_{M,j}^{2}}\right)^{i}\\ &=\frac{1-\left(1-\frac{\sigma_{j}^{2}}{d_{M,j}^{2}}\right)^{k}}{\frac{\sigma_{j}^{2}}{d_{M,j}^{2}}}.\end{split}$
With this we may write $x^{(k)}_{PL}$ as
$\begin{split}x^{(k)}_{PL}&=\sum_{j=1}^{n}\hat{d}_{j}^{k-1}\frac{\sigma_{j}}{d_{M,j}^{2}}\left(u_{A,j}^{T}b\right)v_{M,j}\\ &=\sum_{j=1}^{n}\left(1-\left(1-\frac{\sigma_{j}^{2}}{d_{M,j}^{2}}\right)^{k}\right)\frac{u_{A,j}^{T}b}{\sigma_{j}}v_{M,j}.\end{split}$
@@
    The orthogonality assumption between the right singular vectors of $A$ and $M$ is used in all results herein and may be (approximately) reasonable if the precision of $M$ relative to $A$ is close enough when $M^{T}M$ represents a low precision approximation to $A^{T}A+\alpha^{2}I$.
@@ 4.1 Reformulated Iterative Refinement on the Tikhonov problem
Before proceeding, we consider the general situation that will arise when expanding the $k^{th}$ solution of IR with respect to the solutions of preconditioned Landweber. That is, we consider the simplification of
(21) $\alpha^{2}(M^{T}M)^{-1}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}x^{(q)}_{PL},\quad\begin{cases}i=0,1,\dots,k-1\\ q=0,1,\dots,k-1\end{cases}.$
We first note that $\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}$ may be decomposed using the definition of the preconditioner $M$, the SVD of $A$, and the orthgonality assumption between the right singular vectors of $A$ and of $M$:
$\begin{split}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}&=\left(I-V_{A}\Sigma_{A}^{2}V_{A}^{T}V_{M}D_{M}^{-1}V_{M}^{T}\right)^{i}\\ &=V_{A}\left(I-\tilde{D}_{M}\right)^{i}V_{M}^{T}\end{split}$
where $\tilde{D}_{M}=\Sigma_{A}^{2}D_{M}^{-1}=\text{diag}\left(\sigma_{A,1}^{2}/(\sigma_{M,1}^{2}+\alpha^{2}),\dots,\sigma_{A,n}^{2}/(\sigma_{M,n}^{2}+\alpha^{2})\right)$. With this, we may rewrite (21) following the relation given by (17)
$\alpha^{2}(M^{T}M)^{-1}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}x^{(q)}_{PL}$
$=\alpha^{2}V_{M}D_{M}^{-1}(I-\tilde{D}_{M})^{i}D_{M}^{-1}\hat{D}^{k-1}D_{M}^{-1}\Sigma_{A}^{T}U_{A}^{T}b$
$=\sum_{j=1}^{n}\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{k}\right)\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}.$
@@
$\phi^{(2)}_{j}=\psi^{(2)}_{j}-\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{0}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{1}\right)$
@@
$\phi_{j}^{(k)}=\psi_{j}^{(k)}-\left(\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)\sum_{i=0}^{k-1}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\phi_{j}^{(k-1-i)},\quad j=1,2,\dots,n$
@@
We comment that the filter factors of IR on the Tikhonov problem computed in exact arithmetic are exactly that of Tikhonov regularization. This is stated in Corollary 4.4 below without proof as the argument follows from a straight forward mathematical induction argument.
@@
$\phi^{(k)}_{j}=\frac{\sigma^{2}_{j}}{\sigma^{2}_{j}+\alpha^{2}},\quad\text{for}\,\,\,j=1,2,\dots,n.$

### v2

@@ Abstract
This study investigates the iterative refinement method applied to the solution of linear discrete inverse problems by considering its application to the Tikhonov problem in mixed precision. Previous works on mixed precision iterative refinement methods for the solution of symmetric positive definite linear systems and least-squares problems have shown regularization to be a key requirement when computing low precision factorizations. For problems that are naturally severely ill-posed, we formulate the iterates of iterative refinement in mixed precision as a filtered solution using the preconditioned Landweber method with a Tikhonov-type preconditioner. Through numerical examples simulating various mixed precision choices, we showcase the filtering properties of the method and the achievement of comparable accuracy compared to results computed in double precision as well as another approximate method.
@@
mixed precision, inverse problems, iterative refinement, Tikhonov regularization, preconditioned iterative methods, Landweber
@@ 1 Introduction
In response to advancements in GPU computing architectures that natively support mixed precision computing, Higham and Pranesh in [19] consider the solution of symmetric positive definite (SPD) linear systems and least-squares problems in multiple precisions using iterative refinement (IR). They found that definiteness of SPD matrices could be lost when rounded to a precision with fewer representation bits, which in turn, could result in the failure of a Cholesky factorization when adapting GMRES-based IR strategies; see [2, 3]. It was determined that performing a diagonal shifting scheme could preserve the definiteness of the matrix - which may be viewed as a form of regularization. Our interests in this work are focused on applying IR to least-squares problems whose matrix $A$ is naturally severely ill-conditioned without any assumed underlying structure, to which Tikhonov regularization is often applied. Other works by Higham and others concerning the use of IR to solve linear systems using mixed precision include [4, 13, 20, 24].
@@ 2.1 Filtered Solution of Linear Discrete Ill-posed Problems
Many inverse problems whose approximate solution is given by (1) have been well studied using the singular value decomposition (SVD) of $A=U\Sigma V^{T}$, where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ are orthogonal matrices whose columns represent the left and right singular vectors of $A$, respectively. The diagonal matrix $\Sigma\in\mathbb{R}^{m\times n}$ contains the singular values given by $\sigma_{j}$ for $j=1,2,\dots,n$ on its main diagonal ordered in non-increasing fashion. Using the SVD, one can write the least-squares solution of (1) as a linear combination of the right singular vectors with the $j^{th}$ coefficient given by $u_{j}^{T}b/\sigma_{j}$. Since the singular values of $A$ cluster near numerical zero, the error $e$ contaminating $b$ is propagated into the computed solution for high index values $j$ corresponding to high frequency information [15, 16].
@@ 2.2 Landweber Iteration
where $\zeta\in\left(0,1/\sigma_{1}^{2}\right]$ is a relaxation term with $\sigma_{1}$ the largest singular value of $A$. Without loss of generality, we choose $\zeta=1$ in this work since $A$ can be scaled appropriately; see [10] for further discussion on the choice of $\zeta$ for inverse problems. While robust, this classical variant of Landweber is well known to have slow convergence towards a useful solution which can be understood from viewing its $k^{th}$ solution as a filtering method [14]:
(7) $x^{(k)}=\sum_{j=1}^{n}\Big{(}1-(1-\sigma_{j}^{2})^{k}\Big{)}\frac{u_{j}^{T}b}{\sigma_{j}}v_{j}$
@@ 2.3 Iterative Refinement on the Tikhonov Problem
Variants of fixed and mixed precision IR methods are not new - having been investigated consistently since Wilkinson’s first programmed version in 1948 [32]. A historical treatise of the progression of IR for linear problems may be found in [2, 17]. On a generic linear system, IR computes the $k^{th}$ approximate solution, $x^{(k)}$, in three steps: (i) the $(k-1)^{th}$ residual $r^{(k-1)}$ is computed, (ii) the correction equation $Ah^{(k-1)}=r^{(k-1)}$ is solved, and (iii) the current solution is updated $x^{(k)}=x^{(k-1)}+h^{(k-1)}$. The process may be repeated as necessary until a termination criterion is met.
When considering the application of IR to the Tikhonov problem given by
@@
which was first discussed by Riley in [28], Golub in [12] noted that the three step procedure of IR is equivalent to the procedure of iterated Tikhonov regularization whose $k^{th}$ iterate is given by
(9) $\begin{split}x^{(k)}&=x^{(k-1)}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}\left(b-Ax^{(k-1)}\right)\\ &=x^{(k-1)}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}r^{(k-1)}\end{split}$
which is often used as an iterative technique to approximate the solution of (1). When applying IR to the Tikhonov problem (8) the $k^{th}$ residual, $\hat{r}^{(k)}$, may be written as follows
$\begin{split}\hat{r}^{(k)}&=A^{T}b-\left(A^{T}A+\alpha^{2}I\right)x^{(k)}_{IR}\\ &=A^{T}\left(b-Ax^{(k)}_{IR}\right)-\alpha^{2}x^{(k)}_{IR}\end{split}$
so that the $k^{th}$ iterate may be written recursively as
(10) $\begin{split}x^{(k)}_{IR}&=x^{(k-1)}_{IR}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}\hat{r}^{(k-1)}\\ &=x^{(k-1)}_{IR}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(A^{T}A+\alpha^{2}I\right)^{-1}x_{IR}^{(k-1)}.\end{split}$
@@
1  Input: $A\in\mathbb{R}^{m\times n},\,\,b\in\mathbb{R}^{m},\,\,x_{IR}^{(0)}\in\mathbb{R}^{n},\,\,\alpha>0$
2  Output: $x_{IR}^{(k+1)}\in\mathbb{R}^{n}$
@@ 7        $x_{IR}^{(k+1)}=x_{IR}^{(k)}+h^{(k)}$
From (10) we may observe that the term $\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}\hat{r}^{(k-1)}$ contains the correction term of iterated Tikhonov (9) whose filter factors at the $k^{th}$ iteration take the form
@@
which are described in [12] and [21]. The term $\alpha^{2}\left(A^{T}A+\alpha^{2}I\right)^{-1}x_{IR}^{(k-1)}$ from (10) may be interpreted as a correction term that refines the approximation at each iteration to the solution of (8). Differently than discussed in the aforementioned works, we show that the filter factors of IR on the Tikhonov problem can be written with respect to the $k^{th}$ iterate of iterated Tikhonov regularization (9). This framework lends itself to the filter factor analysis in mixed precision where instead of the $k^{th}$ iterate of (9) we utilize the $k^{th}$ iterate of preconditioned Landweber with a Tikhonov-type preconditioner, which we show in the next section can be written as a filtered solution.
@@ 3 Landweber with a Tikhonov-type Preconditioner
(13) $x^{(k+1)}=x^{(k)}+\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax^{(k)}\right).$
@@
Herein, we utilize the $M$ notation to help simplify our notation throughout this work and to consider the situation that $M^{T}M$ represents an approximate Tikhonov-type preconditioner to $A^{T}A+\alpha^{2}I$. Possible approximations we consider in Section 5 include that of (1) $M^{T}M$ as a low precision SVD approximation and (2) when the structure of $A$ allows for the approximation by a second matrix $C$ whose structure allows for fast and robust matrix-vector products. The former will be our focus when considering IR on the Tikhonov problem in mixed precision. The latter case can arise naturally, for example, in image deblurring problems; see Section 5.
@@
To investigate the filter factors of mixed precision IR on the Tikhonov problem we first describe in Proposition 1 the filter factors of preconditioned Landweber with $M^{T}M=\left(A^{T}A+\alpha^{2}I\right)$. We again emphasize that the main use of the $M$ notation is to delineate between the possible change in structure of the preconditioner $M$ or its representation in a lower precision. Throughout, we will use a capital letter subscript, Q, to denote terms that come from a representation of $Q$, e.g., the SVD of $Q$. Additionally, we utilize subscripts $PL$ and $IR$ on vectors to denote solutions from preconditioned Landweber and IR, respectively.
@@
(15) $x^{(k)}_{PL}=\sum_{j=1}^{n}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{k}\right)\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}$
@@
where we recall that $\hat{A}=AM^{-1}$ and $\hat{x}=Mx$. Using our definition of $M=D_{M}V_{M}^{T}$ and the SVD of $A$, we may rewrite $\hat{A}$ as
@@
(18) $\begin{split}x^{(k)}_{PL}&=M^{-1}\hat{x}^{(k)}_{PL}\\ &=V_{M}D_{M}^{-1}\hat{D}^{k-1}D_{M}^{-1}\Sigma_{A}^{T}U_{A}^{T}b\\ &=\sum_{j=1}^{n}\hat{d}_{j}^{k-1}\frac{\sigma_{A,j}}{d_{M,j}^{2}}\left(u_{A,j}^{T}b\right)v_{M,j}\end{split}$
@@
$\hat{d}_{j}^{k-1}=\sum_{i=0}^{k-1}\left(1-\frac{\sigma_{A,j}^{2}}{d_{M,j}^{2}}\right)^{i}=\frac{1-\left(1-\frac{\sigma_{A,j}^{2}}{d_{M,j}^{2}}\right)^{k}}{\frac{\sigma_{A,j}^{2}}{d_{M,j}^{2}}}.$
Additionally, $d_{M,j}=\left(\sigma^{2}_{M,j}+\alpha^{2}\right)^{1/2}\in\mathbb{R}$ denotes the $j^{th}$ entry from $D_{M}$. With this we may write $x^{(k)}_{PL}$ as
$\begin{split}x^{(k)}_{PL}&=\sum_{j=1}^{n}\hat{d}_{j}^{k-1}\frac{\sigma_{A,j}}{d_{M,j}^{2}}\left(u_{A,j}^{T}b\right)v_{M,j}\\ &=\sum_{j=1}^{n}\left(1-\left(1-\frac{\sigma_{A,j}^{2}}{d_{M,j}^{2}}\right)^{k}\right)\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}.\end{split}$
@@
    The orthogonality assumption between the right singular vectors of $A$ and $M$ is used in all results herein and may be (approximately) reasonable if the precision of $M$ relative to $A$ is close enough when $M^{T}M$ represents a low precision approximation to $A^{T}A+\alpha^{2}I$. To this end, in our numerical experiments in Section 5 we show that the filter factors behave in an approximate Tikhonov-type manner. Additionally, due to comment 1 above, we expect it to be less critical for the filter factors to be exact.
@@ 4.1 Reformulated Iterative Refinement on the Tikhonov problem
Before proceeding, we consider the general situation that will arise when expanding the $k^{th}$ solution of IR with respect to the solutions of preconditioned Landweber, i.e., we consider the simplification of
(22) $\alpha^{2}(M^{T}M)^{-1}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}x^{(k)}_{PL}$
that arises from the ability to rewrite $x^{(k-1-i)}_{IR}$ from the summation term in (20) in terms of components of preconditioned Landweber. We first note that
$\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}$
may be decomposed using the definition of the preconditioner $M$, the SVD of $A$, and the orthgonality assumption between the right singular vectors of $A$ and of $M$:
$\begin{split}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}&=\left(I-V_{A}\Sigma_{A}^{2}V_{A}^{T}V_{M}D_{M}^{-2}V_{M}^{T}\right)^{i}\\ &=V_{A}\left(I-\tilde{D}_{M}\right)^{i}V_{M}^{T}\end{split}$
where $\tilde{D}_{M}=\Sigma_{A}^{2}D_{M}^{-2}=\text{diag}\left(\sigma_{A,1}^{2}/(\sigma_{M,1}^{2}+\alpha^{2}),\dots,\sigma_{A,n}^{2}/(\sigma_{M,n}^{2}+\alpha^{2})\right)$. With this, we may rewrite (22) utilizing the first equality given in (18) so that
$\alpha^{2}(M^{T}M)^{-1}$ $\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}x^{(k)}_{PL}$
$=\alpha^{2}V_{M}D_{M}^{-2}(I-\tilde{D}_{M})^{i}D_{M}^{-2}\hat{D}^{k-1}D_{M}^{-2}\Sigma_{A}^{T}U_{A}^{T}b.$
In conjunction with the second equality of (18) we use that the $j^{th}$ entry of $\alpha^{2}D_{M}^{-2}(I-\tilde{D_{M}})^{i}$ may be expressed as $\frac{\alpha^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\left(1-\frac{\sigma^{2}_{A,j}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{i}$ to write (22) as
$\sum_{j=1}^{n}\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{k}\right)\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}.$
@@
$\begin{split}\phi^{(2)}_{j}&=\psi^{(2)}_{j}-\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{0}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{1}\right)\\ &={\color[rgb]{0,0,0}\psi^{(2)}_{j}-\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{1}\right)}\\ \end{split}$
@@
(23) $\phi_{j}^{(k)}=\psi_{j}^{(k)}-\left(\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)\sum_{i=0}^{k-1}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\phi_{j}^{(k-1-i)},\quad j=1,2,\dots,n$
@@
We comment that the filter factors of IR on the Tikhonov problem computed in exact arithmetic are exactly those of Tikhonov regularization. This is stated in Corollary 4.4 below without proof as the argument follows from a straight forward mathematical induction argument.
@@
(24) $\phi^{(k)}_{j}=\frac{\sigma^{2}_{j}}{\sigma^{2}_{j}+\alpha^{2}},\quad\text{for}\,\,\,j=1,2,\dots,n.$